OpenStudy (anonymous):

Work out the value of each of this expressions, when p=2, q=-3, r=5: p^4+2r^2/pq

OpenStudy (mimi_x3):

(2)^4 + 2(5)^2/(2)(-3) = -11

OpenStudy (anonymous):

how did you get minus eleven, would you mind giving me a quick step by step run through?

OpenStudy (mimi_x3):

typed it in the calculator xD

OpenStudy (anonymous):

LOOOOOOOOOOL

OpenStudy (anonymous):

that cannot be right 2*5^2/-6 is not an integer

OpenStudy (anonymous):

2^4=16 5^2=25 so it is 16+2*25/-6=16-25/3=48/3-25/3=23/3

OpenStudy (mimi_x3):

oh woops my bad i got it wrong xD

OpenStudy (anonymous):

too complicated :-)

OpenStudy (anonymous):

Mimi you could check what I posted, please

OpenStudy (mimi_x3):

LOL, yeah i thought it was easy , it looked easy

OpenStudy (mimi_x3):

okay i will check, Andras (:

OpenStudy (anonymous):

but the answer is - 11

OpenStudy (anonymous):

than u used the brackets incorrectly

OpenStudy (anonymous):

(p^4+2r^2)/pq=-11

OpenStudy (mimi_x3):

see~ i was right, its not that complicated

OpenStudy (anonymous):

andras your method has baffled me

OpenStudy (anonymous):

andras your method has baffled me

OpenStudy (mimi_x3):

chrissy you still need belp ?

OpenStudy (anonymous):

People not use pen and paper any more...:-)

OpenStudy (mimi_x3):

lols , why ? xD

OpenStudy (anonymous):

yes i do

OpenStudy (mimi_x3):

lols, what do you need help with ?

OpenStudy (anonymous):

2^4=16 2(5)^2= 50????

OpenStudy (mimi_x3):

10^2 = 100 sorry for ignoring you before

OpenStudy (anonymous):

so 2(5)^2 is 100

OpenStudy (anonymous):

wat?

OpenStudy (mimi_x3):

yup :D am i wrong, estudier ?

OpenStudy (anonymous):

Exponents first...

OpenStudy (anonymous):

so it is 50 then?

OpenStudy (anonymous):

Yes

OpenStudy (mimi_x3):

oh yeah my bad sorry

OpenStudy (anonymous):

IT's -11 because p^4+2r^2/pq = 16+50/-6 = -11

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes, Mimi had it right since the beginning (but didn't know why)

OpenStudy (anonymous):

thank you for your patience guys.

OpenStudy (mimi_x3):

lols, because i didn't read the question again thats why

OpenStudy (anonymous):

no problem, but wouldnt it be better if u tried to do ur homework alone ???

OpenStudy (mimi_x3):

lols , she cant do it thats why she's asking for herlp xD

OpenStudy (anonymous):

I think the questions are designed to test order of operations...

OpenStudy (anonymous):

She wont do it, i am not trying to be mean so please excuse me if i offended you, but we are limited only by our imagination

OpenStudy (anonymous):

BODMAS, PEDMAS, whatever it's called...

OpenStudy (mimi_x3):

lols , you dont need BODMAS in that question xD

OpenStudy (anonymous):

2(5)^2?....

OpenStudy (mimi_x3):

lols , is BODMAS required ? my bad

OpenStudy (anonymous):

If u just starting, maybe it is not so obvious....

OpenStudy (mimi_x3):

lols , i haven't heard of that term for a while so it wasnt obvious xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

u happy now, chrissy?

OpenStudy (anonymous):

very